### v1

# A Time-Intensity Aware Pipeline for Generating Late-Stage Breast DCE-MRI using Generative Adversarial Models

Abstract
Contrast-enhancement pattern analysis is critical in breast magnetic resonance imaging (MRI) to distinguish benign from probably malignant tumors. However, contrast-enhanced image acquisitions are time-consuming and very expensive. As an alternative to physical acquisition, this paper proposes a comprehensive pipeline for the generation of accurate long-term (late) contrast-enhanced breast MRI from the early counterpart. The proposed strategy focuses on preserving the contrast agent pattern in the enhanced regions while maintaining visual properties in the entire synthesized images. To that end, a novel loss function that leverages the biological behavior of contrast agent (CA) in tissue, given by the Time-Intensity (TI) enhancement curve, is proposed to optimize a pixel-attention based generative model. In addition, unlike traditional normalization and standardization methods, we developed a new normalization strategy that maintains the contrast enhancement pattern across the image sequences at multiple timestamps. This ensures the prevalence of the CA pattern after image preprocessing, unlike conventional approaches. Furthermore, in order to objectively evaluate the clinical quality of the synthesized images, two metrics are also introduced to measure the differences between the TI curves of enhanced regions of the acquired and synthesized images. The experimental results showed that the proposed strategy generates images that significantly outperform diagnostic quality in contrast-enhanced regions while maintaining the spatial features of the entire image. This results suggest a potential use of synthetic late enhanced images generated via deep learning in clinical scenarios.
keywords:
breast cancer , Early-to-late prediction , generative adversarial network (GAN) , magnetic resonance imaging (MRI) , medical image synthesis.
[…]
[O1]organization=Institución Universitaria Pascual Bravo,city=Medellín, country=Colombia
[…]
[O2]organization=Instituto Tecnológico Metropolitano,city=Medellín, country=Colombia
[…]
[O3]organization=Ayudas Diagnósticas SURA S.A.S.,city=Medellín, country=Colombia
[…]
[O4]organization=Morgridge Institute for Research, Dept. Biostatistics and Medical Informatics, University of Wisconsin—Madison,city=Madison, state=WI, country=USA
{graphicalabstract}[Uncaptioned image]
{highlights}
A novel normalization strategy (TI-norm) for DCE-MRI imaging that preserves the kinetic properties of the TI curve across multiple sequences.
A novel loss function (TI-Loss) based on the time-intensity curve that ensures the prevalence of clinically relevant information in the generated images.
Two novel metrics based on the TI curve to objectively evaluate the performance of generated images in terms of clinically relevant information for interpretation.
1 Introduction
Early diagnosis and treatment of breast cancer can significantly improve effectiveness, leading to higher survival rates and reduced cancer-related deaths [1, 2]. Breast cancer is primarily detected through medical imaging, and Dynamic Contrast-Enhanced Magnetic Resonance Imaging (DCE-MRI) is considered the most sensitive method. DCE-MRI does not use radiation and allows the characterization of the internal structures of the tumor [3, 4]. Acquisition of DCE-MRI involves intravenously administration of a chemical gadolinium-based compound known as contrast agent (CA). Due to the accelerated metabolic activity of malignant tumors, they exhibit a characteristic CA absorption pattern that differs from benign and healthy tissues in both speed and intensity over time. This pattern is highly sensitive and makes tumor detection and characterization easier in comparison to other imaging modalities. For radiological findings (i.e. suspicious lesions), the CA absorption pattern is defined by the analysis of the kinetic or time-intensity curve (TI). The TI curve is computed from the variation of the pixel intensity in the location of the maximum intensity point from early or short-term (1-3 min) and late or long-term (8-10 min) responses. This analysis, along with visualization of the internal structure of tumors, makes DCE-MRI one of the most efficient tools for detecting and characterizing breast cancer. [5].
The acquisition of DCE-MRI varies even from hospital to hospital, as typically 5 to 8 screens are taken after CA administration, which usually takes around 15-20 minutes. However, no right number of shots or right time intervals have been established or standardized yet. Furthermore, the large number of images per patient requires a large infrastructure in terms of storage and transmission devices in the database center and the picture archiving and communication system (PACS). Consequently, radiologists must have a long interpretation time to reach a diagnostic conclusion, given the amount of information per patient to analyze. Thus, DCE-MRI becomes extremely restrictive in the number of patients that can be diagnosed; therefore, it is indicated only to individuals at higher risk.
Alternatively, ultra-fast DCE-MRI (UF-DCE-MRI) has emerged to capture kinetic information in the very early post-contrast period by increasing the temporal resolution drastically, which significantly reduces acquisition time. However, increasing temporal resolution also decreases spatial resolution, leading to inaccuracies in the visualization of internal tissue structures [6]. This is why UF-DCE-MRI is not used today as a standalone examination. As another option, abbreviated MRI protocols have been proposed to reduce the number of image sequences that must be acquired, stored, and interpreted [7, 8]. Typically, these protocols include only one early post-contrast image sequence, which allows detection of image findings, but limits tumor characterization because they do not capture the dynamic TI pattern [9]. However, the late response is relevant to distinguish among enhanced masses with benign-appearing morphological features. For example, biologically aggressive cancers will exhibit strong angiogenic activity with fast wash-in and wash-out, whereas fibroadenomas often observe fast wash-in, although wash-out is scarce [10].
On the other hand, deep learning and machine learning models have shown the potential to generate synthetic screens from multiple medical imaging modalities, and several models have been proposed to generate missing or additional image sequences [11, 12, 13]. Some of them have been evaluated as an alternative for the reduction or elimination of CA doses [14, 15]. In the case of breast DCE-MRI, proposed models have focused mainly on synthesizing the early response from the pre-contrast image, which approximates the removal of CA administration [16]. To do so, adaptations of state-of-the-art models [17, 18] and the inclusion of optimization criteria based on the morphological characteristics of the images or tumors [19, 20] have been evaluated. In line with this trend, an intuitive path to obtain a full and useful virtual CA response for diagnosis will imply the synthesis of the late response. However, to our knowledge, no efforts have been reported to address the missing late CA response of breast DCE-MRI, besides our own presented in [21], where we introduced a temporal learning function to exploit the behavior of the TI curve to efficiently generate coherent DCE-MRI.
Since the CA pattern of the TI curve is the most relevant information to distinguish benign from malignant lesions in DCE-MRI, the synthesis of contrast enhancement images must accurately simulate this pattern rather than synthesize a realistic image only. Taking this into account, this paper presents a comprehensive pipeline for the generation of long-term (late) contrast-enhanced breast MRI from its short-term (early) counterpart, which attempts to replicate the behavior of the TI curve of the enhanced regions while maintaining the visual properties in the entire image.
To achieve this, specific components involved in each processing step were addressed. In the same way as our previous work [21], we proposed a new loss function named the Time Intensity Loss function (TI-loss) that takes advantage of the temporal behavior of the CA to guide the training of a generative model to synthesize images that preserve both the CA pattern and spatial features of the generated image. In addition, unlike traditional normalization and standardization methods, we developed a novel strategy that maintains the contrast enhancement pattern across several image sequences at various timestamps. Finally, since the quality of the synthesis process depends on the ability of the model to predict the CA pattern, we propose an evaluation based on two new metrics, the CA pattern score ($\mathcal{CP_{s}}$) that determines the validity of the annotated regions according to their CA pattern (plateau, persistent, and washout), and the average difference in the enhancement ($\mathcal{ED}$), which quantifies the difference between the real and generated enhancement in unannotated regions within the tissue. Pipeline performance was evaluated using a public DCE-MRI image dataset that included studies from 3T and 1.5T scanners with different imaging techniques [22].
The experimental results show that the proposed pipeline generates images that accurately synthesize the contrast enhancement response in terms of the TI curve for both radiological findings (regions of interest - ROI) and unannotated regions, while maintaining visual properties comparable to the real late contrast enhanced images. This result, achieved primarily by implementing the TI-Loss during generative model training, is extremely important, as the evaluation of the baselines indicates that optimizing for pixel-wise quality does not guarantee the prevalence of clinically relevant information. However, the use of TI-Loss alone does not achieve optimal results in terms of clinical performance. The role of the TI-norm was evaluated, demonstrating the prevalence of clinical interpretation and behavior among images to facilitate learning of their relationships while maintaining visual quality in the images. The comparative analysis of the models shows that the proposed pipeline allows the generation of images that preserve spatial properties in terms of pixel metrics, while significantly outperforming them in the clinical interpretation.
Our pipeline presents multiple contributions.
  • A novel normalization strategy (TI-norm) for DCE-MRI imaging that preserves the kinetic properties of the TI curve across multiple sequences.
  • A novel loss function (TI-Loss) based on the time-intensity curve that ensures the prevalence of clinically relevant information in the generated images.
  • Two novel metrics based on the TI curve to objectively evaluate the performance of generated images in terms of clinically relevant information for interpretation.
2 The Time-Intensity curve as diagnostic tool in DCE-MRI
DCE-MRI involves the use of a chemical gadolinium-based compound known as contrast agent (CA). The CA is injected intravenously before the MRI scans to distinguish certain tissue structures. This is possible because tumors or abnormal cells have faster metabolic activity, which makes CA absorb quickly. As a result of the active interaction between the gadolinium and the magnetic field of the MRI scanner, regions containing high concentrations of CA appear as luminous zones in the images. Due to the fast metabolic rate of tumor cells, CA enables precise visualization of the internal tumor composition, size, and shape. Furthermore, it is common to perform a temporal analysis of suspicious lesions across multiple MRI sequences acquired at specific time intervals to determine the most appropriate diagnosis and course of treatment for each patient. These benefits make this technique one of the most effective for cancer detection [5], and the entire collection of sequences and images is usually called DCE-MRI.
The variation pattern in pixel intensity for a zone in a DCE-MRI image over multiple time intervals is typically known as the kinetic or time-intensity (TI) curve. The TI curve is often used to determine the type of lesions based on the contrast enhancement pattern because it is easy to use and effective for both diagnosis and prognosis [23, 24, 25]. From the TI curve, lesions can be classified as probably benign or malignant based on the variation in CA absorption relative to normal tissue in the early and late phases [26]. As illustrated in Figure 1, the TI curve is typified according to its behavior as follows:
  • TI curve type I: This is defined as a progressive or increasing enhancement pattern. It is typical in the case of benign findings, where constant CA absorption but slow evacuation occurs. This increase is often greater than 10% w.r.t. early response over the long run.
  • TI curve type II: This is defined as a plateau pattern. This tendency implies early CA absorption that remains almost constant in the late stage and is equivocal about lesion benignity or malignancy. This variation is within 10% of the late response w.r.t. the early one.
  • TI curve type III: This is defined as a decreasing or washout pattern. This is typical in malignant tumors, where a significant increase in pixel intensities in the short time and a pronounced decrease in the longer time happen. Commonly, the late response decreases lower than 10% with respect to the early counterpart.
Figure 1: Illustration of different TI curves types for DCE-MRI. D0, D1, D2 and D3 stand for the dynamic sequences corresponding to weighted fat-saturated T1. D0 corresponds to pre-contrast, D1 to early post-contrast and D2-D3 to late post-contrast. Blue dotted line shows a persistent TI curve (enhancement increase higher than +10%), orange dashed line shows a plateau TI curve (enhancement change between +10% and -10%) and red dot-dashed line shows a wach-out TI curve (enhancement decrease lower than -10%).
The time-intensity curve has been reported to be an important tool for the clinical evaluation of radiological findings or abnormalities in the case of multimodal DCE MRI. This provides a solid evaluation framework from the pixel analysis level, with reproducible results [27, 28]. The TI curve is widely used in multiple scenarios, as it provides diagnostic value to differentiate benign lesions[29, 30], as well as their type of characterization [31]. In addition, the time intensity curve has been explored and reported as a powerful tool for multiple purposes. To predict the response to treatment in patients [32, 33]. As a key tool for improving and boosting the diagnosis of ultra-fast DCE-MRI by leveraging the increase in temporal resolution [34, 35]. Finally, as an assessment tool to determine the validity of low- or reduced-CA-dose approaches in DCE-MRI [36, 37].
3 Methods
A graphical representation of the proposed pipeline is presented in Figure 2. It is composed by three main stages: Generative model training, late post-contrast synthesis and model evaluation. Before training the generative model, training images are normalized by using a novel normalization technique named time-intensity normalization (TI-norm), that attempts to avoid losing the intensity changes among dynamic sequence, i.e. pre-contrast ant post-contrast images. Then, a Pixel Attention Network (PAN) is used as backbone to synthesize the late post-contrast image from the early counterpart. A novel loss function called TI-loss is here proposed to guide the learning process to minimize the difference among the real and late CA responses. The trained PAN using the TI-loss (TI-PAN) is then used for generating a synthetic late post-contrast image from the TI normalized early image. Finally, the performance of the late contrast synthesis is twofold evaluated. First, conventional image quality metrics (MAE, SSIM and PSNR) between real and generated late post-contrast images are computed to determine the spatial quality of the generated images. Second, two metrics to determine the difference between the real and generated time-intensity curves, allowing the evaluation of the ability to maintain the clinical significance of the CA pattern in the synthesized image.
Figure 2: Graphical representation of our proposed pipeline. It is composed of three main stages: Generative model training, late post-contrast generation and model evaluation. For pre-processing, we proposed the Time Intensity normalization (TI-norm) to ensure the prevalence of the CA behavior among sequences. In the train stage, we proposed the Time Intensity Loss (TI-Loss) to leverage the contrast-enhancement behavior to outperform the diagnostic value of the generated images. The generation stage allows the synthesis of post-contrast images, and for the evaluation stage we proposed a set of metrics based on the TI curve.
3.1 Time Intensity Pixel Attention Network: TI-PAN
As mentioned above, the proposed pipeline is based on synthesizing in the best way the CA pattern. We used an attention-based model called Pixel attention network (PAN) [38]. This model uses attention layers under unique scheme blocks known as self-calibrated convolution (SC-PA) for feature extraction and nonlinear mapping, and nearest-neighbor upsampling (U-PA) for reconstruction. These blocks are considerably more effective than their conventional residual/dense counterparts at the lowest parameter cost. Once the model receives an early-response input image $x_{e}$, the PAN extracts features at the spatial level and stacks multiple SC-PA blocks to generate powerful representations. Subsequently, U-PA blocks are used as reconstruction modules to upsample the features. After reconstruction, a global skip connection path is used in which a bilinear interpolation to the input is performed. Then, the output corresponding to the late response $y_{l^{\prime}}$ is obtained. Finally, to encourage the realism in the image, we added a discriminator network and included an adversarial approach such as in generative adversarial networks (GANs) 3. This also included the pixel-wise loss based on the $L1$ norm to enforce spatial dependency. The full objective for the PAN model is shown in equation 9
$\begin{array}[]{ll}\mathcal{L}_{GAN}(G,D)=&\mathbb{E}_{X_{e},Y_{l}}[logD(x_{e},y_{l})]+\\ &\mathbb{E}_{X_{e}}[log(1-D(x_{e},G(x_{e}))]\end{array}$ (3)
[…]
Time-Intensity Loss for learning late CA response
To help generative models understand contrast agent behavior in tissue, we aim to formulate a general function that models the behavior of the CA response in the images. To achieve this, we created a loss term to determine the difference in the enhanced areas by the contrast agent called the Contrast Enhancement Loss (CELoss) [21]. Specifically, we consider the real early ($x_{e}$) and late ($y_{l}$) responses sampled from data $X_{e}$ and $Y_{l}$ respectively; we compute the real contrast enhanced map ($CE_{r}(x_{e},y_{l})$) by subtracting early from late response (equation 5) using the $L1$ norm. In a similar sense, the contrast enhanced map is also computed for the generated image $CE_{g}$ (equation 6). The model is trained to minimize the difference between $CE_{r}$ and $CE_{g}$ using the $L1$ norm as formulated in equation 7.
$CE_{r}(x_{e},y_{l})=\mathbb{E}_{X_{e},Y_{l}}\left[\ \left\|y_{l}-x_{e}\right\|_{1}\ \right]$ (5)
$CE_{g}(x_{e})=\mathbb{E}_{X_{e}}\left[\ \left\|G(x_{e})-x_{e}\right\|_{1}\ \right]$ (6)
$\mathcal{L}_{CE}(x_{e},y_{l})=\left[\ \left\|CE_{r}(x_{e},y_{l})-CE_{g}(x_{e})\right\|_{1}\ \right]$ (7)
Despite its nice performance and convergence, this method is susceptible to the existing class imbalance between the amount of background and biological signal pixels. This causes gradient diminishing and suboptimal solutions, especially when the $CE$ value is reduced after a few training epochs. To overcome this limitation, we introduce a subtle but important change by replacing the $L1$ with a more flexible function: the Huber loss [39]. The Huber loss uses $L1$ and $L2$ in the estimate to handle outliers, preserves smooth differentiability, and balances accuracy and generalization. It behaves as a quadratic function ($L2$) for values below $\lambda$ and as a linear function ($L1$) for greater values. The formulation of the Huber loss, $\mathcal{L}_{TI}(x_{e},y_{l})$, applied to the contrast enhancement problem is shown in equation 8, where $CE_{r}$ and $CE_{g}$ are functions of $(x_{e},y_{l})$ and $(x_{e})$ respectively, but simplified in notation. We rename this function as the Time-Intensity Loss (TI-Loss):
$\mathcal{L}_{TI}=\\ \left\{\begin{array}[]{ll}1/2\cdot[CE_{r}-CE_{g}]^{2},\ when\ \|CE_{r}-CE_{g}\|\leq\delta\\ \delta\cdot[\|CE_{r}-CE_{g}\|-1/2\cdot\delta]\ otherwise.\end{array}\right.$ (8)
[…]
Hence, combining the TI-Loss with the adversarial approach of the PAN could be defined as in equation 9, where $\lambda$ is a weighting hyperparameter for the loss function.
3.2 Image normalization based on the Time-Intensity pattern
Image normalization plays an important role in image synthesis. Most related works use conventional normalization techniques such min-max and z-score normalization. However, normalizing pixel intensity based on single image statistics can distort diagnostic information because it varies temporally with other precontrast and postcontrast sequences. For this reason, we designed a normalization strategy to ensure that the information in the TI curve is preserved across image sequences. This technique is based on the Z-score standardization with respect to a different reference. For a given precontrast image $x_{p}$, both $\mu$ and $\sigma$ are computed as in the Z-score and the normalized image $x_{p_{TI}}$ is calculated. However, the early ($x_{e}$), real late ($y_{l}$) and generated late ($y_{l^{\prime}}$) images are standardized using the $\mu$ and $\sigma$ values from $x_{p}$. This will ensure that pixel values for all the postcontrast images will reflect the variation in terms of intensity with respect to the precontrast to effectively preserve the contrast enhancement pattern. We named this as Time-Intensity nomralization (TI-norm). To compute the TI-norm of $x_{p},x_{e},y_{l}$ and $y_{l^{\prime}}$, we use equations in 13
$\begin{array}[]{lll}x_{p_{TI}}=\frac{x_{p}-\mu_{x_{p}}}{\sigma_{x_{p}}},&x_{e_{TI}}=\frac{x_{e}-\mu_{x_{p}}}{\sigma_{x_{p}}},&y_{l_{TI}}=\frac{y_{l}-\mu_{x_{p}}}{\sigma_{x_{p}}},\\ \\ &y_{l^{\prime}_{TI}}=\frac{y_{l^{\prime}}-\mu_{x_{p}}}{\sigma_{x_{p}}}\end{array}$ (13)
3.3 Quality evaluation based on Time-Intensity curve
Since the quality of our proposed approach depends on the ability of replicate the expected attributes of the time-intensity curve, we proposed two metrics based on its behavior, besides conventional pixel-based metrics. For a set of precontrast ($x_{p}$), early ($x_{e}$), late ($y_{l}$) and generated ($y_{l^{\prime}}$) responses, a real and a synthetic time-intensity curve is computed and three groups of metrics are estimated:
3.3.1 CA pattern score in annotated regions ($\mathcal{CP}_{s}$)
This metric aims to determine if the CA pattern (plateau, persistent and wash-out) in a region of the image with a small field of view is the same as expected. This approximation aligns with the medical practice, where the TI curve is computed in a single or small valley of pixels to analyze its behavior across multiple image sequences. For a set of regions in all image modalities ($Rx_{p}$, $Rx_{e}$, $Ry_{l}$ or $Rx_{p}$, $Rx_{e}$, $Ry_{l^{\prime}}$) the time intensity curve is computed. To determine the CA pattern for each of them, we calculate the percentage of enhancement ($\mathcal{E}_{ps}$) of the late response w.r.t. the early phase as shown in equation 16. Then, the type of TI curve is estimated following the standard rules shown in equation 15. Finally, the overall multi-class F1 score between them is computed and reported.
$\mathcal{E}_{ps}(Rx_{e},Ry_{l})=\frac{Ry_{l}-Rx_{e}}{Rx_{e}}\times 100(\%)$ (14)
$\mathcal{CP}_{s}=\begin{array}[]{ll}Persistent,&\mathcal{E}_{ps}>+10\%\\ Plateau,&-10\%\leq\mathcal{E}_{ps}\leq+10\%\\ Washout,&\mathcal{E}_{ps}<-10\%\end{array}$ (15)
3.3.2 Average difference of $\mathcal{E}_{ps}$ in regions ($\mathcal{ED}$)
The purpose of this metric is to estimate the difference among both real and generated TI responses within the tissue under different conditions. In this case, the $\mathcal{E}_{ps}$ is computed in pairs of real and generated regions to determine how different the expected score is synthesized. This difference is calculated instead of the CA pattern given its clinical meaning, that is bounded to clinically relevant findings. Formally, given a set of annotated ROIs ($Rx_{p}$, $Rx_{e}$, $Ry_{l}$ or $Rx_{p}$, $Rx_{e}$, $Ry_{l^{\prime}}$) or a set of unannotated regions ($Px_{p}$, $Px_{e}$, $Py_{l}$ or $Px_{p}$, $Px_{e}$, $Py_{l^{\prime}}$) in all image modalities, the TI curve is calculated, as well as the percentage of enhancement ($\mathcal{E}_{ps}$). Then, the $L1$ norm is used to calculate the distance between the real and generated TI curves at late phase from the annotated ROIs ($\mathcal{ED}_{R}$) or unannotated regions ($\mathcal{ED}_{UR}$). The formulation of the score for both scores is shown in equation 16.
$\begin{array}[]{c}\mathcal{ED}_{R}({Rx_{e},Ry_{l},Ry_{l^{\prime}}})=\left\|{\mathcal{E}_{ps}(Rx_{e},Ry_{l})-\mathcal{E}_{ps}(Rx_{e},Ry_{l^{\prime}})}\right\|\\ \mathcal{ED}_{UR}({Px_{e},Py_{l},Py_{l^{\prime}}})=\left\|{\mathcal{E}_{ps}(Px_{e},Py_{l})-\mathcal{E}_{ps}(Px_{e},Py_{l^{\prime}})}\right\|\\ \end{array}$ (16)
3.3.3 Image quality metrics
The purpose of this set of metrics is to evaluate the quality of the images in terms of spatial distribution. For this, we computed the Mean Absolute Error (MAE) to estimate the pixel-to-pixel change among real and generated sequences via spectral fidelity, the Structural Similarity Index Measure (SSIM) to estimate the structural quality of visual perception from three perspectives: correlation loss, luminance distortion, and contrast distortion, and the Peak Signal-to-Noise Ratio (PSNR) to determine the ratio between the maximum possible value (power) of the images and the power of distorting noise that affects their quality. Finally, to better understand the spatial differences in terms of intensity among the real and generated images, we visualized the absolute difference among each pixel intensity value as a heatmap.
4 Results and discussion
4.1 Dataset
To perform all experiments we used the Duke dataset [22]. This dataset includes 922 preoperative MRI imaging sequences of invasive breast cancer patients’ (precontrast and 3 or 4 postcontrast). The patients were between 21 and 89 years old and comprised seven racial and ethnic groups. Tumor size, nuclear grade, and hormone receptor status varied per patient. Ten scanner manufacturers were used to capture MRI images, and patients received three kinds of contrast agents. These constraints maintain heterogeneous parameters and prevent the data from being limited to one MRI configuration. Additionally, expert radiologists annotated at least one region with lesions or tissue anomalies per patient.
Our framework was built using precontrast, the earliest postcontrast (2mins after CA administration), and the latest sequence (6mins after CA administration). In our experiments, 1.5T and 3T images were used independently to train models due to their different visual features in terms of contrast enhancement. The 1.5T and 3T subsets had 393 and 389 patients, respectively. Due to sparse tissue contrast agent response, full DICOM volumes were not used in both cases. A frame-wise selection was performed on the basis of the available annotated ROIs, plus previous and posterior images. Thus, for training and validation, $9,196$ and $1,043$ images were extracted from the 3T images, and $8,941$ and $1,012$ from 1.5T. Additionally, images were normalized using TI-norm, as described in section 3.2.
4.2 Experimental setup
Regarding the PAN model, we ran three tests using the same training configuration and compiled the findings. For the baseline, we adopted the original implementation consisting of 16 SC-PA blocks and convolutions with a kernel size of 40 for downsampling and 24 for upsampling. Additionally, the discriminator network consists of a $1\times 1$ Markovian discriminator [40] with convolutional blocks (Conv2d + InstanceNorm + LeakyReLU) following the architecture as in [41]. Besides, using the same configuration, we trained the model including either CELoss (CE-PAN) or TILoss (TI-PAN) and compiled results to quantify the impact of the losses on the models’ performance.
To perform a comparative analysis considering the CA behavior, we evaluated the quality of the generated images at three different levels: full image, annotated ROIs and unannotated regions. The analysis of the full image is based on conventional pixel-wise metrics to check their realism as other similar works. On the other hand, the evaluation of ROIs annotated is based on the quality of the CA response generated using the metrics described in Section . As the annotations correspond to large portions of tissue, the entire region was not considered. Instead, a small set of pixels was taken ($3\times 3$), where its center corresponds to the maximum intensity point within the ROI.
However, since expected behavior in annotated ROIs could bias the analysis in terms of variability, we also analyzed the generated CA response in tissue regions without annotated clinical findings. For this, regions within the tissue portion of the organs were manually selected in $x_{e}$, and their locations were projected in the rest of the modalities to compile $Px_{p}$, $Px_{e}$, $Py_{l}$, $Py_{l^{\prime}}$. From each region, the TI curve and the increase in enhancement are computed as described in section 4.2. Additionally, to increase the number of samples to analyze, 3 unannotated regions are extracted from each image in the 3T test set, to complete $3,129$ unannotated regions. At this point, the type of TI curve was not considered, as this characterization is usually assigned to lesions or clinically relevant findings. Instead, we expect no significant difference between the generated and expected increase in the enhancement.
4.3 Clinical Performance based on TI curve metrics
Here, we evaluate the accuracy of of the generated images for clinical applications. To achieve this, we used our proposed pipeline to predict the late response images given the early response. Then, the diagnostic value and performance in terms of the TI curve for the annotated ROIs and unannotated regions is analyzed using the aforementioned metrics. Figure 3 a) shows the results for the model predicting the time-intensity response using TI-PAN with the 3T images, where patients are splitted according to their real CA pattern (persistent, plateau, and wash-out). The results show that the estimated CA pattern score for the annotated ROIs is $0.904$ and the expected TI response matches the real response for most patients, with small variations among them.
Additionally, in the most prominent error cases, the proposed pipeline avoids the over-estimation of the late CA response, which is important to avert false positive and false negative cases. Besides, in an in-depth analysis, only 5 cases were miss-estimated for the CA pattern. However, in all these cases the average difference between the real and generated enhancement was ±$\pm 2.5\%$, and the real enhancement was nearby the threshold that determines the type of TI curve (±$\sim\pm 10\%$). Also, all cases were either real persistent or wash-out that were estimated as plateau, which indicates that they are not conclusive and require further examination. These findings are relevant for our purpose since they demonstrates the ability of the model to replicate the expected response in the TI curve for the diagnosis reliably, while preserving the most important information for the clinical analysis and avoiding false positive and false negative cases.
Figure 3: Results for TI curve estimation using our proposed pipeline. Figure a) shows the per-patient TI curve, where black lines and blue points correspond to the reference for the TI curve (pre-contrast and early post-contrast), orange points are the real CA response (late) and green points are the generated response. Additionally, the red dashed line shows the error distance between real and generated enhancement. Patients are organized according to their expected CA pattern (persistent, plateau and wash-out) and their early post-contrast enhancement in ascendingly. Figure b) shows the distribution of the enhancement increase/decrease among real and generated unannotated regions. Besides, dots in the tails of the distributions represent the outliers in terms of the change in increase.
Finally, we aim to determine if there is a significant difference in the CA response between the real and generated group of regions according to their enhancement increase by using the non-parametric Wilcoxon signed rank test, where we chose a significance value of $0.05$. In the test, if the $p$-value rejects the null hypothesis, that means, there is a significant difference between the real and generated CA responses. Otherwise, if the $p$-value fails the test, it means that there is no evidence to support a difference between real and generated CA responses. For the annotated ROIs reported in the figure a), it can be observed that the $p$-values for the persistent, plateau and wash-out TI curves are not able to reject the null hypothesis in any case, which supports that the model is able to generate the CA response following the same real pattern with no significant difference.
However, the analysis on the real annotated ROIs alone exhibits a limited number of samples and cases to analyze, which might bias the evaluation of the quality of the synthesized CA response in terms of generalization. For this reason, we extended the evaluation using unannotated regions within the tissue zones to evidence the changes. Unlike the previous analysis, the determination of the kind of TI curve is not made at this stage due to its clinical significance, which is only done for relevant findings. Instead, the analysis is performed in terms of the expected increase in enhancement to avoid misinterpretations. Plot b) in figure 3 shows the distribution of the enhancement variation for the unannotated real and generated regions and it’s level of statistical significance using the aforementioned test. It can be observed that our pipeline is able to replicate the distribution of the increase in enhancement efficiently, with small differences around the mean. This is also demonstrated in the $p$-value obtained through the test, which indicates no evidence to support a significant difference in the increase in enhancement. Also, it worth notice that outliers for both distributions are located below $-0.25\%$ increase and over $+0.60\%$ increase. These thresholds are important since they indicate that the most susceptible cases of analysis correspond to regions where higher or lower enhancement increase must be analyzed by clinicians, which correlates correctly with the characterization of wash-out and persistent TI curves. For the within-the-distribution the cases, the results show that our pipeline is able to replicate the expected increase in enhancement with small variations.
4.4 Impact of image normalization in the TI-Curve and CA pattern
At this point, we trained TI-PAN using Min-Max, Z-score and TI-norm and generated the synthetic images. In figure 4, first row of a) shows the generated images from all models where no major visual difference is evident. However, to assess the impact of the TI-norm in the diagnostic quality of the images we computed the time-intensity curve of each annotated set of samples $Rx_{p},Rx_{e}$ $Ry_{l}$ and $Ry_{l^{\prime}}$. Using the RadiAnt software [42], we computed the TI curve for the images without any normalization and after normalization. The figure 4 displays time intensity plots in the second row, where it can be seen that the TI curve is destroyed when conventional Min-Max and Z-score are used, in comparison to the non-normalized one. Nevertheless, the TI-PAN model still tries to mimic the given behavior when using Min-Max or Z-score with certain similarity. This is a critical issue for the diagnosis since the information of the curve itself is degraded significantly, and it distorts its interpretability afterwards despite the high visual correlation among real and generated late post-contrast. Nevertheless, our proposed time-intensity normalization effectively preserves the CA behaviour in the real and generated images, which approaches the generated TI curve to a real clinical interpretation.
Figure 4: Impact of normalization in the diagnostic information. a) Top row displays the early (input) and late (output) responses, besides the generated images using the TI-PAN model. Bottom row shows the time intensity plots for annotated enhanced ROI, where blue continuous line corresponds to the computed real time-intensity curve in late and after normalization, and orange dashed line is the generated time-intensity curve by the model. b) Shows a comparative plot of image quality (PSNR) and CE difference using the model and different normalization strategies in annotated ROIs and unannotated regions.
We also extended the evaluation by comparing the predicted contrast enhancement difference and the image quality in annotated and unannotated 3T images, and results are shown in b). Although in this case the Z-score norm obtains best performance in terms of image quality, it falls very short in terms of the expected estimation of the CA pattern, achieving suboptimal performance. Similar behavior is also obtained using Min-Max norm, where pixel error is lower but diagnostic quality remains suboptimal. In both cases, the TI-PAN model shows a decrease in performance and exhibits a critical limitation in the clinical scenario, despite using TI-Loss. In the case of TI-norm, it guarantees the prevalence of the TI curve to be leveraged by the TI-Loss during training. These features allow to obtain best scores in the estimation of the CA pattern while achieving similar image-based quality (25.4 in PSNR). This makes TI-norm an important component in generating high-quality diagnostic information in an image synthesis setting, which has not been reported before in the medical imaging synthesis field.
4.5 Quantitative and Qualitative evaluation of synthesized Images
An evaluation in the visual quality of all the generated images at spatial feature level was performed. Figure 5 shows some generated sample results for images in the test set using all of the evaluated models together. Additionally, as images in the test set had at least an annotated ROI, they have been also compared visually. Finally, the difference heatmaps are also shown. In the case of 1.5T images, the generated samples show enough realism in the visual domain for all of the models, with exception of certain tissue textures and details. This is coherent with the previous clinical result where narrow margin of difference was obtained in the visual domain. However, in case of the 3T images, a smaller difference in the real and generated is evidenced when using the TILoss, especially in the annotated region where the difference in the brighter zone is lower. Besides, table 1 also reports the MAE, SSIM and PSNR for all models at full image and ROI leves, where no single model obtains the best scores consistently. In contrast, no large margin is evidenced among models, and slight differences are found. This result is in line with other cutting-edge studies that have shown good results in creating useful images from post-contrast sequences [18, 11].
Figure 5: Comparative plots for visual image quality. Heatmaps represents the difference of the real vs the generated images in the same scale. Annotated ROIs are also shown in augmented projection. Best seen zoomed in.
4.6 Models Comparison
Finally, we evaluated the PAN model performance and compare it with the CE-PAN and TI-PAN using the scores described in section 4.2. Table 1 shows the comparative performance between all models in 1.5T and 3T images.
In the case of 1.5T images, pixels metrics are very similar, and best performances are divided for PAN (SSIM full image, PSNR and SSIM ROIs) and TI-PAN (MAE and PSNR full image). These results align correctly with the qualitative images shown previously. Similarly, in the case of the 3T images, CE-PAN obtains the best performance in most metrics, except PSNR in ROIs. Nevertheless, the difference margin between metrics is not large, and most models achieve similar values, where the largest difference is obtained in the PSNR with 0.3 and 0.1 for 3T and 1.5T respectively. However, the case of the TI metrics is different, since TI-PAN obtains the best scores compared to PAN and CE-PAN. In the case of annotated ROIs extracted from tissue, the TI-PAN increases the estimation of CE intensity by $0.05\%$ in the case of 1.5T and $0.03\%$ in the case of 3T. Although the increase is not large, it might be crucial in the diagnosis of patients since it might represent the difference between the different types of curves within the tissue.
A similar behavior is obtained for the unannotated regions, where an increase in CE intensity estimation is obtained for 1.5T images ($0.11\%$) and 3T images ($0.07\%$). This result is relevant because it shows the ability of our model to estimate the intensity of CE, regardless of whether an anomaly is present in the tissue or not. This also shows that the TI-PAN model does not rely on the presence of tumor, but leverages the biological behavior of the CA in tissue. Finally, a large margin is reported in the $E_{ps}$ score, where an increase of $0.20\%$ for both 1.5T and 3T images. The increase in performace is caused by the TI-Norm and TI-Loss altogether, since they exploit the TI curve to synthesize the images.
Duke 1.5T Duke 3T
PAN CE-PAN TI-PAN PAN CE-PAN TI-PAN
Pixel (images) MAE 0.555 0.554 0.543 0.706 0.701 0.736
PSNR 27.05 26.96 27.08 25.64 25.71 25.54
SSIM 0.689 0.688 0.687 0.691 0.696 0.688
Pixel (ROI) MAE 0.701 0.7 0.713 0.924 0.922 0.938
PSNR 21.69 21.66 21.58 21.22 21.23 21.24
SSIM 0.514 0.511 0.509 0.582 0.584 0.579
TI metrics $\mathcal{ED}_{UR}$ 0.611 0.612 0.493 0.346 0.404 0.334
$\mathcal{ED}_{R}$ 0.174 0.178 0.122 0.164 0.17 0.14
$\mathcal{CP}_{s}$ 0.629 0.621 0.827 0.62 0.614 0.818
Table 1: Comparative results among all models using the test sets. MAE, SSIM and PSNR are reported as spatial feature metric for real/generated pair of images in the test set at two levels: full images and ROIs. Additionally, the metrics based in the TI curve ($\mathcal{ED}_{UR}$, $\mathcal{ED}_{R}$ and $\mathcal{CP}_{s}$) are also reported.
5 Discussion
Late enhancement, which is performed 8 or 10 min after injection of the contrast agent, provides relevant information to distinguish benign from probably malignant lesions in a Breast MRI study. However, acquiring it has the drawback of additional acquisition time and costs. Abbreviated protocols have gained attention as a cost-effective alternative in the medical field. However, concerns about their accuracy persist due to the omission of the late contrast enhancement phase, which, as mentioned, is important for an accurate diagnosis. Therefore, the development of alternatives for estimating the late contrast enhancement response is a valuable tool for integrating abbreviated protocols into clinical practice.
In this work, we propose a pipeline that uses a deep learning-based generative model to predict the late response of a DCE-MRI from its early counterpart, which is optimized to preserve the expected biological behavior of the contrast agent by introducing the Time-Intensity loss function. The pipeline leverages the behavior of contrast agents at multiple levels, demonstrating its ability to reliably outperform the synthesis of clinically relevant information while preserving the visual properties of the images compared to existing models.
To our knowledge, there are no previous works that have addressed the problem of the synthesis or generation of late-enhanced images for breast DCE-MRI. However, some works that attempt to generate contrast-enhanced images have been reported. Those works have shown adequate performance to generate realistic images with acceptable performances from a low-level feature point of view. However, no evaluation of the ability to reproduce the behavior of the contrast agent pattern has been considered, which is the key factor in determining the feasibility of using it in clinical practice. The results presented in this work demonstrate that image quality metrics are not completely suitable to optimize clinical information beyond visual similarity. Instead, task-specific types of functions, such as the proposed TI-Loss, could help achieve clinically interpretable results.
The synthesizer performance was evaluated two ways: using stablished pixel-wise metrics, and evaluating the ability to replicate clinically relevant information (TI curve). Experimental results for the pixel-wise metrics show enough realism and similar performance in the full image setting, and also in annotated ROIs and unannotated regions. For the second group, we formulate two new metrics named the CA pattern score ($\mathcal{CP}_{s}$) and the average difference in the percentage of enhancement ($\mathcal{ED}$). In this case, the TI-PAN results significantly outperform other models by a large margin in both metrics. These results are relevant since they demonstrate that our pipeline is able to maintain image quality during the synthesis process while outperforming the clinical performance of the model and preserving the full interpretability of the diagnostic information.
Image normalization is a standard stage in any image synthesis process. We evaluated multiple normalization techniques commonly used in the field of medical image synthesis and found that the most common normalization approaches (i.e. Min-max and Z-score) significantly modify the behavior of the TI curve, destroying its objective interpretation, as radiologists do in their daily practice. For this reason, we introduced the time-intensity normalization (TI-norm) strategy that preserves the diagnostic information required for physician interpretation. Our results in enhanced selected regions show that the proposed TI-norm preserves the information in the TI curve beyond the visual cues.
This work differs from others because we proposed an entire pipeline where all stages were carefully intervened in to preserve important information for the clinical scenario, rather than only a single component, a model or function to process the data. Experimental results are promising for the implementation of late contrast-enhanced synthesis in a clinical setting. However, it must be evaluated by a controlled clinical study to determine the validity of the diagnostic information in a real assessment, particularly under different acquisition variables, variability among clinicians, inter-reader interpretation, etc. In particular, performance with different contrast agents and dosages must be evaluated, considering that the pipeline is grounded in the qualification of the CA response and the TI-curve. These variations cause differences in visualization and might alter the training process. This kind of study is out of the scope of this work and will be evaluated as a future line of research.
6 Conclusion
In this work, we introduced a strategy for predicting late-contrast-enhanced images of its early counterpart in breast DCE-MRI. We have proven the discrepancy between image quality during the synthesis and the prevalence of clinically significant information determined by contrast enhancement behavior. Based on this concept, we introduce three main innovations for clinical image synthesis: 1) a novel loss function that leverages the biological behavior of the contrast agent (CA) in tissue to optimize generative models, 2) a normalization strategy based on the time-intensity (TI) curve that preserves the diagnostic information required for interpretation and diagnosis, and 3) two performance evaluation metrics that measure how synthesized late enhanced images are able to replicate the TI-curve in enhanced regions. The proposed strategy demonstrates suitability for image synthesis by preserving the image quality of models but significantly outperforming the interpretability of the TI curve, which showed promising potential in clinical applications.
Funding
This work was supported by Minciencias, Colombia under the “programa de becas de excelencia doctoral del bicentenario”, BPIN 2021000100029.